### v1

@@ 8 Conclusion
The application of Transformers in the intrusion detection domain represents a significant advancement due to their powerful Attention mechanisms and ability to handle diverse data types. Attention-based IDS leverages the capability of Transformers to focus on relevant parts of the input data, enhancing the detection of subtle anomalies in complex network traffic. CNN/LSTM-Transformer-based IDS combines the spatial feature extraction strengths of CNNs and the temporal sequence modeling capabilities of LSTM with Transformers’ Attention mechanisms, providing a robust solution for capturing both spatial and temporal anomalies in data.
ViT-based IDS adapts the principles of vision Transformers, originally designed for image processing, to the IDS context. This approach allows for sophisticated pattern recognition in visual representations of network traffic or system logs, improving the detection accuracy of complex attacks. GAN-Transformer-based IDS integrates GANs with Transformers, where GANs generate synthetic data to augment the training set, helping the Transformer model to learn better and generalize from limited intrusion data. This combination enhances the system’s ability to detect rare or novel attack patterns. Furthermore, FL-Transformer-based IDS addresses data privacy concerns by enabling the training of Transformer models across decentralized data sources without needing to centralize sensitive data. This approach not only preserves privacy but also benefits from diverse data, improving the robustness and accuracy of IDS models. Overall, the integration of Transformers with various ML paradigms and architectures offers a versatile and powerful toolkit for developing advanced IDS solutions, capable of addressing the ever-evolving landscape of cybersecurity threats.
The application of LLMs for IDS demonstrates significant potential in enhancing security across various environments. GPT-based IDS and BERT-Transformer-based IDS variants, such as DeepLog, LogRobust, HitAnomaly, LogSy, LogAnomaly, NeuralLog, LogBERT, and LAnoBERT, leverage DL to identify complex patterns and anomalies in data. These models excel in processing vast amounts of data from diverse sources, including network traffic, IoT device logs, and ICSs. Their ability to learn from contextual information makes them highly effective for anomaly detection. However, challenges such as data heterogeneity, real-time processing requirements, and model interpretability remain. Addressing these challenges requires careful integration with existing security infrastructures, domain-specific knowledge for data preprocessing, and ongoing model maintenance. Despite these challenges, LLMs hold promise for advancing IDS capabilities, offering robust and adaptable solutions for detecting and mitigating cyber threats.

### v2

@@ 8 Conclusion
The field of intrusion detection is rapidly evolving, driven by advancements in Transformer-based models. These models, with their powerful Attention mechanisms, have outperformed traditional detection systems, offering enhanced capabilities in identifying subtle anomalies within complex network environments. However, several challenges still limit their full potential, including issues related to data diversity, real-time processing, computational resources, and model interpretability. The survey underscores the growing importance of integrating Transformers into IDSs to improve cybersecurity, particularly by capturing both spatial and temporal anomalies through models incorporating CNN/LSTM layers. These approaches reduce false positive rates and enhance detection accuracy. This survey also explores novel Transformer-based methods, such as ViT-based and GAN-Transformer-based IDSs, which offer advanced capabilities in detecting complex attack patterns and improving model performance through synthetic data generation, showcasing the versatility of Transformers in strengthening IDSs.
Moreover, the relevance of Transformers extends beyond conventional IDSs. Transformer-based models, such as those employing LLMs for intrusion detection, show significant potential in addressing challenges associated with heterogeneous data sources, and advanced attack vectors. By utilizing context-aware models like GPT-based or BERT-based IDSs, these systems offer improved adaptability in complex cybersecurity environments. More over, this survey includes a dedicated section on real-world case studies of Transformers and LLMs in IDSs. This section examines the reviewed literature, focusing on empirical case studies for Transformer-based IDSs and LLM-based IDSs. By assessing the application of these models in real-world IDS environments, the section highlights the practical effectiveness of these algorithms, their deployment challenges, and provides valuable insights for future replication and extension in operational settings.
Despite the challenges outlined in this survey, both Transformers and LLMs offer immense potential for advancing the field of intrusion detection. As research in this domain continues to evolve, it is anticipated that these models will play a critical role in enhancing the security, privacy, and resilience of digital infrastructures. Future innovations in Transformer and LLM-based IDSs, alongside advancements in data handling, real-time processing, and model interpretability, promise to revolutionize intrusion detection, ensuring these systems can effectively tackle the ever-evolving cybersecurity threats.